# A Survey of Optimization-based Task and Motion Planning: From Classical To Learning Approaches

### v2

@@ I Introduction
In recent years, robotic systems are rapidly transitioning from structured factory floors to unstructured human-centric environments. To this end, the demand continues to grow for a planning system that enables robots to efficiently perform complex, long-horizon tasks, as exemplified in Fig. I. To achieve this level of autonomy, robots must be capable of generating and executing feasible and efficient motion plans that allow them to interact with their environment and complete assigned tasks. This complex problem is often framed as robot task and motion planning (TAMP), which breaks a complex, often intractable planning problem into a hybrid symbolic search and a set of local motion planning problems, where each sub-problem is tractable to solve.
@@ I-A Survey Goals and Roadmap
Q1 and Q2 motivate us to explore the key components and critical features of optimization-based TAMP, including problem formulation (Sec. II), domain representation (Sec. III), task planning (Sec. IV), and motion planning (Sec. V). Q3 seeks to present the current strategies and remaining challenges of optimization-based TAMP (Sec. VI), and inspire improved TAMP frameworks that efficiently manage complex task structures and robot dynamics. A particular focus of the discussion is on the interaction between task planning and motion planning layers. Q4 addresses advancements in learning-based methods with the intent of synthesizing these elements for the enhancement of TAMP frameworks. The discussion on Q4 is interleaved with the classical methods to place the learning approaches into proper context. An overview of the stucture of the survey is illustrated in Fig. I.
@@ II-C Motivating Example
We introduce a tabletop manipulation task as a representative example to illustrate the formulations and algorithms discussed in this survey. As illustrated in Fig II-C, the task involves a robot manipulator, denoted as $R$, and three distinct movable objects labeled as $A,B,C$. The primary objective of this task is to stack the objects such that the final height of object $A$ is maximized. Additionally, the robot should exert a minimal amount of control effort to achieve this task.
@@ III-D Generating Domain Knowledge by LLMs
Problem Description: The problem description is: You have 3 blocks. B is on top of C. C is on top of A. A is on the table. B is clear. Your arm is empty. Your goal is to move the blocks, so that the final height of block A is as high as possible. Provide me with the problem PDDL file that describes the planning problem directly without further explanations. Problem File: (define (problem blocks-problem) (:domain blocks) (:objects A B C - block) (:init (clear B) (on B C) (on C A) (on-table A) (arm-empty) (= (height A) 0) (= (height B) 0) (= (height C) 0) ) (:metric (height A)) )
@@ VI-A1 Search Heuristics in TAMP
Many search heuristics in TAMP attempt to solve a relaxation of the underlying motion planning problem, in order to obtain an estimation of the feasibility and cost of the action. For example, in TAMP for navigation problems [154], Euclidean distance in 2D space serves as an admissible and easily computable heuristic function, which improves planning efficiency while guaranteeing task-level optimality. However, in planning domains with high-dimensional configuration spaces, it is often difficult to generate an analogous distance measure that estimates the action costs. One intuitive approach is to evaluate the action feasibility and cost based on the initial and final configurations of the robot or object while ignoring the intermediate trajectories. For example, inverse kinematics (IK) is commonly used to reason about the feasibility of the initial and final robot poses in the action without generating the full trajectory [163, 123]. For collaborative robot manipulation, the angular displacement of a manipulated object is often used as heuristics [16][164] proposes an object-centric representation of manipulation constraints that unifies TAMP into a single heuristic search that is amenable to existing AI planning heuristics. [8] proposes a multi-stage method to solve the TAMP: 1) optimizing over the final configuration given an action sequence, 2) optimizing over all kinematics configurations at symbolic state transitions, and 3) optimizing over the entire trajectory. The first two stages effectively act as heuristics during tree search to check the geometric feasibility of a given symbolic action sequence, while the costly TO is only conducated in the final stage.
@@ VI-C Learning for combined TAMP
[189]
Feasibility Classifier
object states; motion feasibility
offline
tabletop manipulation
[190]
Feasibility Classifier
abstract action sequence; plan feasibility
online
tabletop manipulation
[27]
Feasibility Classifier
abstract action sequence, goal descriptions, images, and object states; plan feasibility
offline
mobile manipulation
[38]
State and Action Abstractions
low-level transitions with object states and actions; symbolic predicates
offline
mobile manipulation
[59]
State and Action Abstractions
low-level transitions with images, segmentation masks, and actions; symbolic predicates
offline
mobile manipulation
[87]
State and Action Abstractions
low-level transitions with images and actions; skill feasibility
offline + online
tabletop manipulation
[88]
State and Action Abstractions
low-level transitions with object states and actions; skill execution cost
offline + online
tabletop manipulation
[32]
Control Policy
low-level transitions with object states and actions; rewards
online
tabletop manipulation
[191]
Control Policy
low-level transitions with object states and actions
offline
tabletop manipulation
[192]
Control Policy; State and Action Abstractions
low-level transitions with object states and actions; symbolic predicates
offline
tabletop manipulation
[193]
Combinatorial Decisions
MICP solutions
offline
free-flyer & dexterous grasping
@@ VII Future Challenges and Opportunities
TAMP for Locomotion and Manipulation: Although the duality between locomotion and manipulation [236] means that they can be viewed as equivalent problems, the distinct nature of locomotion and manipulation presents challenges in applying TAMP more complex robotic systems, e.g., humanoid robots with loco-manipulation capabilities. From the TAMP perspective, both involve hybrid planning for dynamic contact interactions between the robot and the environment but differ in the representation and frequency of the hybrid events. Manipulation tasks are usually object-centric [2] and occur at lower frequencies for contact switching, focusing on precise interactions with objects. In contrast, locomotion involves robot-centric motions with higher frequency contact switching, typically handled through a hierarchical approach of contact planning followed by trajectory generation [23], often facilitated by a scentroidal trajectory planning approach [237, 238]. This difference in representation and operational frequency raises important research questions in developing a unified TAMP framework for loco-manipulation, requiring a balanced integration of these aspects. [18] shows promises in integrating graph search and TO for long-horizon loco-manipulation tasks by handling the object-centric and robot-centric tasks separately and transferring offline generated plans to online execution. [239] illustrates the recent trend to use hierarchical RL to solve whole-body loco-manipulation problems for humanoid robots with dexterous hands. Still, TAMP for unified locomotion and manipulation remains a challenge, due to complicating factors such as the complexity of dexterous grasping and locomotion planning in uneven terrains.

### v3

@@ I Introduction
In recent years, robotic systems are rapidly transitioning from structured factory floors to unstructured human-centric environments. To this end, the demand continues to grow for a planning system that enables robots to efficiently perform complex, long-horizon tasks, as exemplified in Fig. 1. To achieve this level of autonomy, robots must be capable of generating and executing feasible and efficient motion plans that allow them to interact with their environment and complete assigned tasks. This complex problem is often framed as robot task and motion planning (TAMP), which breaks a complex, often intractable planning problem into a hybrid symbolic search and a set of local motion planning problems, where each sub-problem is tractable to solve.
@@ I-A Survey Goals and Roadmap
Q1 and Q2 motivate us to explore the key components and critical features of optimization-based TAMP, including problem formulation (Sec. II), domain representation (Sec. III), task planning (Sec. IV), and motion planning (Sec. V). Q3 seeks to present the current strategies and remaining challenges of optimization-based TAMP (Sec. VI), and inspire improved TAMP frameworks that efficiently manage complex task structures and robot dynamics. A particular focus of the discussion is on the interaction between task planning and motion planning layers. Q4 addresses advancements in learning-based methods with the intent of synthesizing these elements for the enhancement of TAMP frameworks. The discussion on Q4 is interleaved with the classical methods to place the learning approaches into proper context. An overview of the structure of the survey is illustrated in Fig. 2.
@@ II-C Motivating Example
We introduce a tabletop manipulation task as a representative example to illustrate the formulations and algorithms discussed in this survey. As illustrated in Fig 3, the task involves a robot manipulator, denoted as $R$, and three distinct movable objects labeled as $A,B,C$. The primary objective of this task is to stack the objects such that the final height of object $A$ is maximized. Additionally, the robot should exert a minimal amount of control effort to achieve this task.
@@ III-D Generating Domain Knowledge by LLMs
Problem Description: The problem description is: You have 3 blocks. B is on top of C. C is on top of A. A is on the table. B is clear. Your arm is empty. Your goal is to move the blocks, so that the final height of block A is as high as possible. Provide me with the problem PDDL file that describes the planning problem directly without further explanations. Problem File:
(define (problem blocks-problem)
(:domain blocks) (:objects A B C - block) (:init (clear B) (on B C) (on C A) (on-table A) (arm-empty) (= (height A) 0) (= (height B) 0) (= (height C) 0) ) (:metric (height A)) )
@@ VI-A1 Search Heuristics in TAMP
Many search heuristics in TAMP attempt to solve a relaxation of the underlying motion planning problem, in order to obtain an estimation of the feasibility and cost of the action. For example, in TAMP for navigation problems [154], Euclidean distance in 2D space serves as an admissible and easily computable heuristic function, which improves planning efficiency while guaranteeing task-level optimality. However, in planning domains with high-dimensional configuration spaces, it is often difficult to generate an analogous distance measure that estimates the action costs. One intuitive approach is to evaluate the action feasibility and cost based on the initial and final configurations of the robot or object while ignoring the intermediate trajectories. For example, inverse kinematics (IK) is commonly used to reason about the feasibility of the initial and final robot poses in the action without generating the full trajectory [163, 123]. For collaborative robot manipulation, the angular displacement of a manipulated object is often used as heuristics [16][164] proposes an object-centric representation of manipulation constraints that unifies TAMP into a single heuristic search that is amenable to existing AI planning heuristics. [8] proposes a multi-stage method to solve the TAMP: 1) optimizing over the final configuration given an action sequence, 2) optimizing over all kinematics configurations at symbolic state transitions, and 3) optimizing over the entire trajectory. The first two stages effectively act as heuristics during tree search to check the geometric feasibility of a given symbolic action sequence, while the costly TO is only conducted in the final stage.
@@ VI-C Learning for combined TAMP
[189] Feasibility Classifier object states; motion feasibility offline tabletop manipulation
[190] Feasibility Classifier abstract action sequence; plan feasibility online tabletop manipulation
[27] Feasibility Classifier abstract action sequence, goal descriptions, images, and object states; plan feasibility offline mobile manipulation
[38] State and Action Abstractions low-level transitions with object states and actions; symbolic predicates offline mobile manipulation
[59] State and Action Abstractions low-level transitions with images, segmentation masks, and actions; symbolic predicates offline mobile manipulation
[87] State and Action Abstractions low-level transitions with images and actions; skill feasibility offline + online tabletop manipulation
[88] State and Action Abstractions low-level transitions with object states and actions; skill execution cost offline + online tabletop manipulation
[32] Control Policy low-level transitions with object states and actions; rewards online tabletop manipulation
[191] Control Policy low-level transitions with object states and actions offline tabletop manipulation
[192] Control Policy; State and Action Abstractions low-level transitions with object states and actions; symbolic predicates offline tabletop manipulation
[193] Combinatorial Decisions MICP solutions offline free-flyer & dexterous grasping
@@ VII Future Challenges and Opportunities
TAMP for Locomotion and Manipulation: Although the duality between locomotion and manipulation [236] means that they can be viewed as equivalent problems, the distinct nature of locomotion and manipulation presents challenges in applying TAMP more complex robotic systems, e.g., humanoid robots with loco-manipulation capabilities. From the TAMP perspective, both involve hybrid planning for dynamic contact interactions between the robot and the environment but differ in the representation and frequency of the hybrid events. Manipulation tasks are usually object-centric [2] and occur at lower frequencies for contact switching, focusing on precise interactions with objects. In contrast, locomotion involves robot-centric motions with higher frequency contact switching, typically handled through a hierarchical approach of contact planning followed by trajectory generation [23], often facilitated by a centroidal trajectory planning approach [237, 238]. This difference in representation and operational frequency raises important research questions in developing a unified TAMP framework for loco-manipulation, requiring a balanced integration of these aspects. [18] shows promises in integrating graph search and TO for long-horizon loco-manipulation tasks by handling the object-centric and robot-centric tasks separately and transferring offline generated plans to online execution. [239] illustrates the recent trend to use hierarchical RL to solve whole-body loco-manipulation problems for humanoid robots with dexterous hands. Still, TAMP for unified locomotion and manipulation remains a challenge, due to complicating factors such as the complexity of dexterous grasping and locomotion planning in uneven terrains.